LASER INDUCED FLUORESCENCE SPECTROSCOPY OF JET COOLED SiCN : ROTATIONAL ANAL-YSIS OF THE HOT BANDS

\underline{MASARU FUKUSHIMA} and TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

We have generated SiCN in supersonic free jet expansions, and observed the laser induced fluorescence ( LIF ) of the vibrationally hot bands of the Ã ^{2}Δ – X̃ ^{2}Π transition. We have measured dispersed fluorescence ( DF ) spectra from the single vibronic levels ( SVL’s ), Ã (01^{1}0) ^{2}Φ and ^{2}Π, and rotationally resolved LIF excitation spectra of the two hot bands, Ã (01^{1}0) ^{2}Φ – X̃ (01^{1}0) ^{2}Δ and Ã (01^{1}0) ^{2}Π – X̃ (01^{1}0) ^{2}Σ^{(−)}. The rotational energy levels were reasonably analyzed as those of the ^{2}K′ – ^{2}K′′ transitions, but their line intensities calculated from the Hönl-London factors derived in the intermediate case between Hund’s case (a) and (b) could not reproduce the observed spectra. The Hönl-London factors derived in the ^{2}Λ′ – ^{2}Λ′′ ( ^{2}Δ – ^{2}Π ) transition reasonably reproduced the spectra. It indicates that coupling between the electronic orbital and vibrational angular momenta is weak in the SiCN ^{2}Δ – ^{2}Π system, and a basis set of |Λ v_{2} l Σ; J P M_{J}⟩, so-called ”l-basis”, better describes the system than that of |Λ v_{2} K Σ; J P M_{J}⟩.